# CATD: Unified Representation Learning for EEG-to-fMRI Cross-Modal Generation

### v1

@@ 5 Conclusion
In this work, a novel CATD framework is proposed for the cross-modal conversion of functional neuroimages, specifically the synthesis of BOLD signals from EEG signals. To fully exploit the high temporal resolution of EEG signals, the DTFS module was designed to increase the sampling rate of the EEG signal as a conditioned signal, achieving temporal resolution enhancement of the synthesized BOLD signal. By constructing the CAB module, the alignment of high-dimensional heterogeneous functional neuroimages in the hidden space was realized. Qualitative and quantitative experimental results demonstrate that the proposed framework effectively achieves cross-modal synthesis from EEG to BOLD signals. The effectiveness of CAB was validated through ablation experiments, and the framework’s value was illustrated in practical application scenarios through medical decision support experiments. Future studies will focus on further optimizing the model and improving the quality of the reconstructed signals to achieve more comprehensive functional neuroimage synthesis.
Acknowledgment

### v3

@@ 5 Conclusion
In this work, a novel CATD framework is proposed for the cross-modal conversion of functional neuroimages, specifically the synthesis of BOLD signals from EEG signals. To fully exploit the high temporal resolution of EEG signals, the DTFS module was designed to increase the sampling rate of the EEG signal as a condition signal, achieving temporal resolution enhancement of the synthesized BOLD signal. By constructing the CAB module, the alignment of high-dimensional heterogeneous functional neuroimages in the hidden space was realized. Qualitative and quantitative experimental results demonstrate that the proposed framework effectively achieves cross-modal synthesis from EEG to BOLD signals. The effectiveness of CAB was validated through ablation experiments, and the framework’s value was illustrated in practical application scenarios through medical decision support experiments. Future studies will focus on further optimizing the model and improving the quality of the generated signals to achieve more comprehensive functional neuroimage synthesis.